1. Support department cost allocation—sequential method

Becker Tabletops has two support departments (Janitorial and Cafeteria) and two production departments (Cutting and Assembly). Relevant details for these departments are as follows:

Support Department Cost Driver
Janitorial Department Square footage to be serviced
Cafeteria Department Number of employees
Janitorial
Department
Cafeteria
Department
Cutting
Department
Assembly
Department
Department costs $310,000    $169,000 $1,504,000    $680,000
Square feet 50 5,000 1,000 4,000
Number of employees 10 3    30 10   

Allocate the support department costs to the production departments using the sequential method.

Allocate the support department with the highest department cost first.

Cutting
Department
Assembly
Department
Janitorial Department $ $
Cafeteria Department $ $

------------------------------------------------------------------------------------------------------

2.

Joint Cost Allocation—Net Realizable Value Method

Lily’s Lemonade Stand makes three types of lemonade: pure, raspberry, and strawberry. The lemonade is produced through a joint mixing process that costs a total of $30 per batch. One batch produces 32 cups of pure lemonade, 21 cups of strawberry lemonade, and 21 cups of raspberry lemonade. After the split-off point, all three lemonades can be sold for $0.80 per cup, but strawberry and raspberry lemonade can be processed further by adding artificial coloring and flavoring and sold for $0.95 and $1.00 per cup, respectively. It is estimated that these additional processing costs are $0.75 and $1.80 per batch for strawberry and raspberry lemonade, respectively.

Allocate the joint costs of production to each product using the net realizable value method.

Joint Product Allocation
Pure lemonade $
Strawberry lemonade
Raspberry lemonade
Totals $

------------------------------------------------------------------------------------------------

3.

Support department cost allocation—reciprocal services method

Becker Tabletops has two support departments ( Janitorial and Cafeteria) and two production departments (Cutting and Assembly). Relevant details for these departments are as follows:

Support Department Cost Driver
Janitorial Department Square footage to be serviced
Cafeteria Department Number of employees
Janitorial
Department
Cafeteria
Department
Cutting
Department
Assembly
Department
Department costs $310,000    $169,000 $1,504,000    $680,000
Square feet 50 5,000 1,000 4,000
Number of employees 10 3    30 10   

Allocate the support department costs to the production departments using the reciprocal services method.

Cutting
Department
Assembly
Department
Janitorial Department cost allocation $ $
Cafeteria Department cost allocation $ $
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Becker Tabletops
Sequential Method Janitorial Cafeteria Cutting Assembly Total
Square Feet        5,000.00            1,000.00            4,000.00         10,000.00
Square Feet % 50.00% 10.00% 40.00% 100.0000%
Number of Employees                 30.00                 10.00                 40.00
Number of Employees % 75.00% 25.00% 100.0000%
Total costs       310,000.00    169,000.00 1,504,000.00       680,000.00 2,663,000.00
Janitorial cost to be allocated- Square Feet %      (310,000.00)    155,000.00         31,000.00       124,000.00                        -  
Total                        -      324,000.00    1,535,000.00       804,000.00 2,663,000.00
Cafeteria cost to be allocated- Number of Employees % (324,000.00)       243,000.00         81,000.00
Total cost after allocation                        -                       -   1,778,000.00       885,000.00 2,663,000.00
Lily’s Lemonade
Workings for NRV A B C=A*B
Calculation of final sales value Cups Sell Price Market Value
Pure lemonade                  32.00                0.80                 25.60
Strawberry lemonade                  21.00                0.95                 19.95
Raspberry lemonade                  21.00                1.00                 21.00
Total                  74.00                 66.55
Calculation of NRV Pure lemonade Strawberry lemonade Raspberry lemonade Total
Market Value                  25.60              19.95                 21.00
Further processing costs                        -                  0.75                   1.80
NRV                  25.60              19.20                 19.20                 64.00
Weight of NRV 40.00% 30.00% 30.00% F
G See F H=F*G
Allocation of Joint costs Joint Cost Weight of base Allocated cost
Pure lemonade                  30.00 40.00%                 12.00
Strawberry lemonade                  30.00 30.00%                   9.00
Raspberry lemonade                  30.00 30.00%                   9.00
Total                 30.00
Becker Tabletops
Reciprocal Method Janitorial Cafeteria Cutting Assembly Total
Square Feet                        -          5,000.00            1,000.00            4,000.00         10,000.00
Number of Employees                  10.00                     -                   30.00                 10.00                 50.00
Square Feet % 0.00% 50.00% 10.00% 40.00% 100.00%
Number of Employees % 20.00% 0.00% 60.00% 20.00% 100.00%
Total costs       310,000.00    169,000.00 1,504,000.00       680,000.00 2,663,000.00
Let cost of Janitorial is $ X and Cafeteria is $ Y
So, X= 310000+ 0.20Y
Y= 169000+ 0.50X
Substituting the value of X in equation 2:
Y = 169000 + 0.50(310000 + 0.20Y)
Y = 169000 + 115000+.010Y
Y – 0.010Y = 115000+169000
0.90Y = 384,000
Y = 384000/0.90
Y = $ 426,667
Substituting the value of Y in equation 1:
So, X= 310000+ .020Y
So, X= 310000+ 0.20*426667
So, X= 310000+ 85333
X= $ 395,333
Allocation Summary Janitorial Cafeteria Cutting Assembly Total
Square Feet %                        -   50.00% 10.00% 40.00% 100.00%
Number of Employees % 20.00%                     -   60.00% 20.00% 100.00%
Department cost       310,000.00    169,000.00 1,504,000.00       680,000.00 2,663,000.00
Total costs
Janitorial      (395,333.00)    197,666.50         39,533.30       158,133.20                        -  
Cafeteria          85,333.00 (426,667.00)       256,000.20         85,333.80                        -  
Cost after allocation                        -      (60,000.50)